OpenStudy (anonymous):

What is the least common denominator of the fraction 7 over 3 x plus the fraction 2 over x plus the fraction x over x minus 1 ? @SolomonZelman

OpenStudy (anonymous):

Help PLEASE

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{7}{3x}+\frac{2}{x}+\frac{x}{x-1} }\)

OpenStudy (solomonzelman):

3x(X-1) would be the common denominator.

OpenStudy (anonymous):

HELP WITH SOME MORE PLEASE

OpenStudy (solomonzelman):

I might not be online, but will see...

OpenStudy (anonymous):

What is the least common denominator of the fraction 3 x over x plus 1 plus the fraction x plus 1 over 2 x plus the fraction 5 over x ?

OpenStudy (anonymous):

What is the least common denominator of the fraction 4 over x plus the fraction 2 x over x minus 1 plus the fraction x plus 3 over 2 x ?

OpenStudy (anonymous):

What is the least common denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x ?

OpenStudy (anonymous):

And that is all

OpenStudy (anonymous):

Are you going to help @Jake33